# One Stone with Two Birds: A Null-Text-Null Frequency-Aware Diffusion Models for Text-Guided Image Inpainting

### v4

@@ 2.3.3 Null-Text Mid-Frequency Aware Denoising Process
We resort to another null-text denoising process guided by the denoised mid-frequency band from the second text-guided denoising process (Sec.2.3.2) for the $t$-th step of the early stage. Particularly, we initialize the denoising process with random Gaussian noise $z^{mid}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and null-text $c_{\emptyset}$ as input to the denoising network $\epsilon_{\theta}(z_{T}^{mid},T,\tau_{\theta}(c_{\emptyset}))$. We further exploit the denoised mid-frequency result from ${z}_{t}^{text}$ , to substitute the denoised mid-frequency band ${z}^{mid}_{t}$ from the last null-text denoising process, throughout a plug-and-play Denoised Mid-Frequency Band layer (as illustrated in Fig.4(b)) to yield $\tilde{z}^{mid}_{t}$, which is formulated below:
$\begin{split}\tilde{z}^{mid}_{t}=&\,\text{IDCT}\left(\text{DCT}({z}_{t}^{text})\odot m_{\text{mid}}+\text{DCT}({z}^{mid}_{t})\odot(1-m_{\text{mid}})\right),\end{split}$ (8)
@@ 2.4 Late Stage of Text-Guided Denoising Process
Following the discussion above, for final late stage of denoising process, we perform the text-guided denoising process based on the early null-text mid-frequency guided denoising process (Sec.2.3.3), the low-frequency band within the masked regions is denoised throughout the path to text prompts alignment, together with the stable mid-frequency band (see Fig.2(b)) from the early text-guided denoising process (Sec.2.3.2), the masked regions are well inpainted. However, the low-frequency band for unmasked regions biases a lot from the ground truth from the early null-text denoising process (Sec.2.3.3), and augmented by the influence from the text prompts for late stage. To this end, we substitute the unmasked regions from early stage of the diffusion process into the denoising results $z^{mid}_{t}$ for the $t$-th step during the late stage to yield $\hat{z}^{mid}_{t}$, which is formulated as:
$\begin{split}\hat{z}^{mid}_{t}=z^{gt}_{T-t}\odot m_{z}+z^{mid}_{t}\odot(1-m_{z}),\end{split}$ (11)
where $z^{gt}_{T-t}$ donates the forward diffused results with the (Tt)(T-t)-th step. Building upon the noisy sample $\hat{z}^{mid}_{t}$, the text prompt $c$, along with the step $t$, we perform the denoising network $\epsilon_{\theta}(\hat{z}^{mid}_{t},t,\tau_{\theta}(c))$ to get the denoising result ${z}^{mid}_{t-1}$ in the latent space for the next (t1)(t-1)-th step. After $\lambda T$ steps, the denoised latent representation ${z}^{mid}_{0}$ is reconstructed into the pixel space via the VAE decoder $\mathcal{D}$ to generate the final image inpainting output. (See the pseudo algorithm in Sec.C of the Appendix)
@@ II. Early Stage
1:Initialize $z_{T}^{un}$, $z_{T}^{text}$, $z_{T}^{mid}\sim\mathcal{N}(0,1)$.
@@
7:  Substitute mid-frequency band of $z_{t}^{mid}$ with the corresponding counterpart of $z_{t}^{text}$ via Eq.8–10;
8:  Compute $z_{t-1}^{mid}$ from $z_{t}^{mid}$, conditioned on $C_{\emptyset}$;
@@ III. Late Stage
2:  Substitute unmasked regions of $z_{t}^{mid}$ with the corresponding counterpart of $z_{t}^{gt}$ via Eq.11;
3:  Compute $z_{t-1}^{mid}$ from $z_{t}^{mid}$, conditioned on $C$;
@@
5:Output the final inpainted image $I^{\prime}=D(z_{0}^{mid})$.

### v5

@@ 2.3.3 Null-Text Mid-Frequency Aware Denoising Process
We resort to another null-text denoising process guided by the denoised mid-frequency band from the second text-guided denoising process (Sec.2.3.2) for the $t$-th step of the early stage. Particularly, we initialize the denoising process with random Gaussian noise $z^{in}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and null-text $c_{\emptyset}$ as input to the denoising network $\epsilon_{\theta}(z_{T}^{in},T,\tau_{\theta}(c_{\emptyset}))$. We further exploit the denoised mid-frequency result from ${z}_{t}^{text}$ , to substitute the denoised mid-frequency band of ${z}^{in}_{t}$ from the last null-text denoising process, throughout a plug-and-play Denoised Mid-Frequency Band layer (as illustrated in Fig.4(b)) to yield $\tilde{z}^{in}_{t}$, which is formulated below:
$\begin{split}\tilde{z}^{in}_{t}=&\,\text{IDCT}\left(\text{DCT}({z}_{t}^{text})\odot m_{\text{mid}}+\text{DCT}({z}^{in}_{t})\odot(1-m_{\text{mid}})\right),\end{split}$ (8)
@@ 2.4 Late Stage of Text-Guided Denoising Process
Following the discussion above, for final late stage of denoising process, we perform the text-guided denoising process based on the early null-text mid-frequency guided denoising process (Sec.2.3.3), the low-frequency band within the masked regions is denoised throughout the path to text prompts alignment, together with the stable mid-frequency band (see Fig.2(b)) from the early text-guided denoising process (Sec.2.3.2), the masked regions are well inpainted. However, the low-frequency band for unmasked regions biases a lot from the ground truth from the early null-text denoising process (Sec.2.3.3), and augmented by the influence from the text prompts for late stage. To this end, we substitute the unmasked regions from early stage of the diffusion process into the denoising results $z^{in}_{t}$ for the $t$-th step during the late stage to yield $\hat{z}^{in}_{t}$, which is formulated as:
$\begin{split}\hat{z}^{in}_{t}=z^{gt}_{T-t}\odot m_{z}+z^{in}_{t}\odot(1-m_{z}),\end{split}$ (11)
where $z^{gt}_{T-t}$ donates the forward diffused results with the (Tt)(T-t)-th step. Building upon the noisy sample $\hat{z}^{in}_{t}$, the text prompt $c$, along with the step $t$, we perform the denoising network $\epsilon_{\theta}(\hat{z}^{in}_{t},t,\tau_{\theta}(c))$ to get the denoising result ${z}^{in}_{t-1}$ in the latent space for the next (t1)(t-1)-th step. After $\lambda T$ steps, the denoised latent representation ${z}^{in}_{0}$ is reconstructed into the pixel space via the VAE decoder $\mathcal{D}$ to generate the final image inpainting output. (See the pseudo algorithm in Sec.C of the Appendix)
@@ II. Early Stage
1:Initialize $z_{T}^{un}$, $z_{T}^{text}$, $z_{T}^{in}\sim\mathcal{N}(0,1)$.
@@
7:  Substitute mid-frequency band of $z_{t}^{in}$ with the corresponding counterpart of $z_{t}^{text}$ via Eq.8–10;
8:  Compute $z_{t-1}^{in}$ from $z_{t}^{in}$, conditioned on $C_{\emptyset}$;
@@ III. Late Stage
2:  Substitute unmasked regions of $z_{t}^{in}$ with the corresponding counterpart of $z_{t}^{gt}$ via Eq.11;
3:  Compute $z_{t-1}^{in}$ from $z_{t}^{in}$, conditioned on $C$;
@@
5:Output the final inpainted image $I^{\prime}=D(z_{0}^{in})$.